OpenStudy (anonymous):

8(2x – 14) + 13 = 4x – 27 Write the steps you will use to solve the equation, and explain each step. What value of x makes the equation true?

OpenStudy (yoloshroom):

Mutiply it all out. So 8*2x+8*-14+13=4x-27 Next 16x-112+13=4x-27 , We Simplify. so 16x-99=4x-27 , then we subtract 4x from both sides. 16x-4x-99=-27 simplify. 12x-99=-27 Next add 99 to both sides. 12x=99-27 simplify , 12x=72 divide both sides by 12 to get X by itself. x=72/12 x=6

OpenStudy (yoloshroom):

that work for you? or do i need more detail

OpenStudy (anonymous):

you make it seem so complicated o_o

OpenStudy (yoloshroom):

I just tried to do it step by step XD , shall i make it easier?

OpenStudy (yoloshroom):

would you prefer of i did?

OpenStudy (anonymous):

yes plz :)

OpenStudy (yoloshroom):

Multiply out 8(2x-14) so we get 16x-112+13=4x-27 move X to one side, and all other terms to the other side :) 16x-4x=112-13-27 then we simplify the equation so 12x=72 then we divide by 12 so x=6

OpenStudy (yoloshroom):

i hope that's better!